A heat engine absorbs 760 kJ heat from a source at 380K. It rejects (1) 650 kJ, (ii) 560 kJ, (iii) 504 kJ of heat to sink at 280 K. State which of these represent a reversible, an irreversible and an impossible cycle.
1
Expert's answer
2015-12-07T12:53:42-0500
Solution
According to 2d law of Thermodynamics:
deltaT/T1 = A/Q
A = deltaT *Q / T1 = 100 * 760 / 380 = 200 kJ.
Rejected: 760 – 200 = 560 kJ – ii (reversible cycle).
Answer: (ii) 560 kJ - reversible cycle.